if Xand Y component of vector A are 4 and 6 metre and the X and Y component of vector A +  vector B are 10 and 9 metre calculate the length of vector B and its angle with respect to x-axis.

Grade:11

1 Answers

Piyush Kumar Behera
7 years ago
A=4i+6j
A+B=10i+9j
so B=6i+3j
angle made with x axis=>tan\theta=y/x=3/6=1/2
\theta=tan-1½
hence B=6i+3j and angle made with x axis=tan-1½